OpenStudy (anonymous):

Factor this polynomial completely: x4– 25.

OpenStudy (anonymous):

(x^2+5)(x^2-5)

OpenStudy (anonymous):

difference of two squares: \[a^2-b^2=(a+b)(a-b)\] here \[a=x^2\] \[b=5\] so get \[(x^2+5)(x^2-5)\]

OpenStudy (anonymous):

\[x^2+5\] does not factor, but \[x^2-5\] does as \[(x+\sqrt{5})(x-\sqrt{5})\]